# TF-SASM: Training-free Spatial-aware Sparse Memory for Multi-object Tracking

### v1

@@ Abstract
Multi-object tracking (MOT) in computer vision remains a significant challenge, requiring precise localization and continuous tracking of multiple objects in video sequences. This task is crucial for various applications, including action recognition and behavior analysis. Key challenges include occlusion, reidentification, tracking fast-moving objects, and handling camera motion artifacts. Past research has explored tracking-by-detection methods and end-to-end models, with recent attention on tracking-by-attention approaches leveraging transformer architectures. The emergence of data sets that emphasize robust reidentification, such as DanceTrack, has highlighted the need for effective solutions. While memory-based approaches have shown promise, they often suffer from high computational complexity and memory usage. We propose a novel sparse memory approach that selectively stores critical features based on object motion and overlapping awareness, aiming to enhance efficiency while minimizing redundancy. Building upon the MOTRv2 model, a hybrid of tracking-by-attention and tracking-by-detection, we introduce a training-free memory designed to bolster reidentification capabilities and preserve the model’s flexibility. Our memory approach achieves significant improvements over MOTRv2 in the DanceTrack test set, demonstrating a gain of 1.1% in HOTA metrics and 2.1% in IDF1 score.
@@ I Introduction
Historically, research efforts in Multiple Object Tracking (MOT) have primarily pursued two distinct directions: tracking-by-detection [sort, deepsort] methodologies and end-to-end models [centertrack, tracktor, memot, memotr]. Tracking-by-detection methods typically rely on object detectors to identify individual objects within each frame, followed by the application of filtering techniques like the Kalman filter to establish correspondence between bounding boxes across the video sequence [sort, ocsort, bytetrack]. Some innovations in this realm have focused on enhancing reidentification by leveraging deep features [deepsort]. Additionally, heuristic approaches, such as camera motion compensation [deepocsort, yi2024ucmctrack], have been proposed to address specific challenges in the tracking domain.
In contrast, end-to-end models represent a more recent paradigm shift, consolidating both detection and association tasks within a unified deep-learning framework. Initially, earlier work often used Convolutional Neural Network (CNN) models for detection, followed by the propagation of detected objects into subsequent frames [tracktor, centertrack]. However, with the introduction of the DETR (DEtection TRansformer) model [detr], which leverages the Transformer architecture [transformer] and stands as a state-of-the-art detector at that time, attention-based tracking approaches [trackformer, transtrack, motr], a subtype of end-to-end models, have emerged, demonstrating promising results in various tracking scenarios.
@@
Our proposed method offers a streamlined solution: a sparse memory that selectively stores critical features based on object motion and overlapping awareness. This approach leverages the intuition that an object’s degree of deformation correlates directly with its movement, thereby minimizing redundancy and enhancing long-temporal Ouinformation.
To validate our hypothesis, we incorporate our memory mechanism into the MOTRv2 [motrv2] model due to its inherent flexibility. MOTRv2 represents a novel hybrid approach, combining tracking-by-attention and tracking-by-detection methodologies, thus harnessing the strengths of robust detectors and a transformer-like architecture for tracking tasks. This architecture exhibits adaptability across diverse scenarios due to its detector and tracker components separation. By extending this flexibility, we introduce a novel training-free memory mechanism for object tracking, thereby enhancing the association performance of the model.
@@ II-A Tracking-by-detection
Previous methods in multi-object tracking (MOT) have typically followed a similar pipeline: initially, a detector is trained to localize objects in each frame independently, followed by implementing a matching strategy to associate these detections across the video sequence.
One prevalent matching strategy is exemplified by the Simple Online and Realtime Tracking (SORT) [sort] algorithm, which employs a Kalman filter to predict the next state of an object and then utilizes bipartite matching, facilitated by the Hungarian algorithm, to associate these predictions with new detections. Building upon SORT, DeepSORT [deepsort] incorporates additional object feature matching, thereby improving performance in reidentification tasks.
Another noteworthy approach is BYTETrack [bytetrack], which acknowledges that low-confidence detections may arise due to occlusions or blur, potentially leading to missed tracking if disregarded entirely. BYTETrack leverages these low-confidence detections in a second-round association, effectively reducing false-negative cases.
OC-SORT [ocsort] introduces an innovative update mechanism for the Kalman filter: when an object is rediscovered after a period of missed tracking, the filter’s parameters are adjusted based on the observation. Deep OC-SORT [deepocsort] extends this concept by incorporating adaptive updates of deep features for object reidentification, guided by the detection confidence scores. Furthermore, Deep OC-SORT introduces a camera motion compensation technique with adjustments tailored to observation-centric updates, enhancing robustness in tracking scenarios influenced by camera movements.
@@ II-B End-to-end Approaches
End-to-end object tracking models primarily utilize two approaches: tracking-by-regression and tracking-by-attention. In tracking-by-regression, these methods aim to propagate object detections from one frame to the next, rather than relying on matching detections between consecutive frames as in the traditional tracking-by-detection paradigm. In this approach, the object detector not only provides frame-wise detections but also replaces the data association. For instance, previous endeavors such as Tracktor [tracktor] employ a regression head to determine whether an object exists in the subsequent frame. Alternatively, in CenterTrack [centertrack], the model predicts detections based on object centers encoded as a heatmap in the preceding frame.
When DETR [detr] was proposed for the object detection problem, it introduced an alternative approach: a single model capable of both detection and tracking using Transformer architecture [transformer]. Following this architectural trend, a recent breakthrough is TrackFormer [trackformer], which pioneers object tracking within an end-to-end framework using an attention mechanism in Transformer architecture. Specifically, TrackFormer treats objects from the previous frame as object queries in the subsequent frame, similar to the object detection process employed in the DETR model [detr]. This paradigm shift enables TrackFormer to effectively track objects while harnessing the inherent capabilities of Transformer-based architectures.
Similarly, MOTR [motr] shares an identical concept with TrackFormer. However, while TrackFormer learns within two adjacent frames, MOTR proposes mechanisms for training with multiple frames by aggregating track queries using transformer architecture. In contrast, TransTrack [transtrack] employs an Intersection over Union (IoU) matching mechanism instead of relying solely on the regression head to determine new object appearances or objects from the previous frame.
II-C Hybrid Approaches
Although tracking-by-attention methods have demonstrated efficacy, they encounter a fundamental challenge in adapting to diverse object classes, necessitating training data for both detection and tracking simultaneously. To tackle this challenge, MOTRv2 [motrv2] introduces a novel approach that combines the strengths of tracking-by-detection and tracking-by-attention methods. This is achieved by leveraging a robust pretrained detector tailored to specific object classes and employing an attention mechanism to propagate objects to subsequent frames via cross-attention mechanisms.
However, the original work lacks the necessary memory mechanisms for robust reidentification tasks. Indeed, several studies have showcased the efficacy of memory mechanisms in enhancing reidentification [memotr, memot] within the context of the Multiple Object Tracking (MOT) problem. Unlike Single Object Tracking, which involves a smaller number of stored objects, MOT presents a challenge due to its larger number of objects. Departing from prevalent memory-based approaches that employ deep networks to select and densely store object features at each frame, we emphasize using a sparse memory to mitigate memory usage, particularly for retaining long temporal information. This system selectively stores crucial object features based on deformation and occlusion awareness.
@@ III Method
Figure 1: Our proposed method seamlessly integrates into the MOTRv2 [motrv2] model as a training-free module, preserving MOTRv2’s flexibility for various applications. As depicted in the figure, the pipeline takes proposals generated by YOLOX [yolox2021] and previously tracked objects (detailed in Section III-B) as input for MOTRv2. The Spatial-aware Sparse Memory (SASM), further explained in Figure 2, then processes MOTRv2’s outputs.
In this section, we present our method, which builds on the MOTRv2 [motrv2] model, renowned for its flexibility and superior tracking capabilities. Specifically, we revisit the MOTRv2 architecture and introduce our memory module designed to enhance reidentification performance. The pipeline depicted in Figure 1 illustrates how our memory module takes the embeddings of the object produced by MOTRv2 in each frame and undergoes a filtering process to retain the most important features of the object. These selected features are then propagated to the next frame to facilitate object localization. The filtering process consists of two modules: Spatial-Aware Sparse Memory (SASM) and Overlapping-Aware Feature Selector (OFS), which we will elaborate on in Sections III-C and III-D, respectively.
III-A Problem definition
In the Multiple Object Tracking (MOT) problem, we are presented with a video $\mathcal{V}$ consisting of frames $I_{t}$. The primary objectives involve detecting objects within each frame and tracking a specified set of object classes $\mathcal{C}$. For each class, the algorithm aims to output a collection of bounding boxes corresponding to these objects in each video frame $\mathcal{V}$, along with their respective identities. These detections form tracklets $\mathcal{T}$, sequences of detections associated with the same object throughout the entire video.
III-B Revisiting MOTRv2
End-to-end multiple-object tracking frameworks often suffer from sub-par detection performance compared to tracking-by-detection approaches that rely on standalone object detectors. To address this issue, the authors propose integrating the YOLOX [yolox2021] object detector to provide detection priors for MOTR [motr].
III-B1 Proposal query generation
In this approach, for each frame $t$, YOLOX generates $N_{t}$ proposals $P$, each characterized by center coordinates $(x_{t},y_{t})$, height $h_{t}$, width $w_{t}$, and confidence score $s_{t}$. Initially, a query mechanism similar to the DAB-DETR [liu2022dabdetr] architecture is employed for the first frame, using object coordinates as reference points and a shared learnable embedding as proposal queries. The confidence scores are also integrated into the proposal queries $\mathbf{q}_{t}$ using sine-cosine positional encoding.
III-B2 Proposal propagation
@@
In subsequent frames, the object queries $\mathbf{q}_{t-1}^{tr}$ and predictions from the previous frame are propagated and concatenated with the proposals from the new frame $P_{t}$. A regression head at the end of the model then determines whether the detected objects are newly appearing or previously tracked objects.
III-B3 Discussion
While object features inherently retain some historical information, prior studies have shown the benefits of incorporating explicit object memory to improve reidentification tasks. To this end, we propose a novel memory integration directly into the MOTRv2 model without requiring fine-tuning. Our experiments reveal that dense memory may struggle to maintain sufficient object information over time in scenarios with a limited number of frames, resulting in redundancy and challenges in matching rapidly evolving object states in subsequent frames. Therefore, we introduce a novel mechanism for selecting object features based on movement and overlapping, enabling the use of smaller memory while preserving longer temporal information.
III-C Spatial-aware Sparse Memory
In this study, we propose a hypothesis linking an object’s deformation to its movement. Accordingly, we utilize the object’s center’s displacement as a metric to determine whether its feature should be stored in memory. The memory pipeline is depicted in Figure 2. The algorithm proceeds through the following steps:
Figure 2: The Spatial-aware Sparse Memory (SASM) takes the output of MOTRv2 as input, which includes both object queries and coordinates. From the object coordinates, the module filters out small-displacement objects to accumulate distance to the next frame, and stores only objects with large displacement, as described in Section III-C. The Overlapping-aware Feature Selector (OFS) plays a crucial role in selecting the best features of each instance since the last memory update, as described in Section III-D.
@@ III-C Spatial-aware Sparse Memory
    Thresholding Moving Distance: If the computed moving distance $d_{i,t}$ exceeds a predefined threshold, indicating significant motion, the object’s features are stored in memory. This step ensures that only objects undergoing substantial movement are considered for memory storage.
@@
    Accumulating Moving Distance: If the moving distance $d_{i,t}$ does not surpass the threshold, it is accumulated with the moving distance of the object in the next frame $t+1$. This accumulation accounts for gradual movement across consecutive frames.
    $d_{i,t+1}=d_{i,t+1}+d_{i,t}$
  4. 4.
@@
These steps collectively form a coherent framework for selective features stored in memory, leveraging object movement as a key criterion for memory allocation and ensuring robustness in object tracking across frames. This sparse memory strategy not only addresses the issue of bias towards the latest appearances of objects in feature representation but also facilitates the capture of longer temporal information when compared with dense memory.
III-D Overlapping-aware Feature Selector
Based on our empirical observations, we have noted that object features are often erroneously stored during occlusion or when objects overlap, introducing noise into the stored features. To address this issue, we propose a method for selectively choosing object features to store in memory.
A simplistic approach might involve waiting until the object is in a less obstructed location; however, our experiments reveal that this strategy leads to significant temporal gaps between the last feature stored in memory and the current frame. Therefore, instead of merely storing the object’s latest query feature, we opt to select the object feature from the frame with the lowest Intersection over Union (IoU) with other objects since the last feature update.
This combination not only minimizes the inclusion of erroneous or redundant features but also ensures a smoother continuity in feature representation across frames. By selecting object features based on their spatial relationships with other objects, we aim to enhance our memory storage mechanism’s discriminative power and robustness.
Mathematically, we define the stored feature $\mathbf{\hat{q}}_{i,t}$ at frame $t$ as follows:
@@ III-D Overlapping-aware Feature Selector
where $\mathbf{q}_{i,t}^{tr}$ represents the feature of object has index $i$ at frame $t$, and $\mathbf{q}_{i,k}^{tr}$ is the feature of object has index $i$ at the frame $k$ with the highest IoU score. This selection process aims to retain more reliable object features in memory, thereby reducing noise and enhancing tracking accuracy.
As a result, this selective feature storage strategy contributes to the overall efficiency and effectiveness of our tracking system, enabling more accurate and reliable object tracking over extended durations. The pseudo-code of the method is represented in detail in Algorithm 1
0:  $\mathbf{q}_{t}^{tr},\mathbf{q}_{t-1}^{tr}$
1:  Set $O$ with object IDs exists in both $\mathbf{q}_{t}^{tr},\mathbf{q}_{t-1}^{tr}$
2:  Compute $\text{IoU}_{i,t}$
3:  Update $\mathbf{\hat{q}}_{i,t}^{tr}$ as Eq. 3
4:  for $i\in O$ do
5:     Calculate $d_{i,t}$ as Eq. 1
6:     if $d_{i,t}>threshold$ then
7:        Update memory by $\mathbf{\hat{q}}_{i,t}^{tr}$
8:        Update $\mathbf{q_{i,t+1}}$ as Eq. 2
9:     else
10:        Accumulate $d_{i,t+1}$ by $d_{i,t}$
11:        $\mathbf{q}_{i,t+1}=\mathbf{q}_{i,t}^{tr}$
12:     end if
13:  end for
14:  return  $\mathbf{q}_{t+1}$
Algorithm 1 Spatial-aware Sparse Memory
@@ IV-A Dataset and Metric
In this paper, we evaluate our methods on the DanceTrack [dancetrack] dataset. This dataset requires tracking multiple objects with uniform appearance and diverse motion. It comprises 100 videos, with 40 for training (annotations public), 25 for validation (annotations public), and 35 for testing (annotations unpublic). To ensure fair comparisons, we assess our methods using well-known metrics as employed in previous works [motr, motrv2]: HOTA [hota] to evaluate our method and analyze the contribution decomposed into detection accuracy (DetA) and association accuracy (AssA); MOTA [clearmot]; IDF1 [id-metrics] metrics.

### v2

@@ Abstract
Multi-object tracking (MOT) in computer vision remains a significant challenge, requiring precise localization and continuous tracking of multiple objects in video sequences. The emergence of data sets that emphasize robust reidentification, such as DanceTrack, has highlighted the need for effective solutions. While memory-based approaches have shown promise, they often suffer from high computational complexity and memory usage due to storing feature at every single frame. In this paper, we propose a novel memory-based approach that selectively stores critical features based on object motion and overlapping awareness, aiming to enhance efficiency while minimizing redundancy. As a result, our method not only store longer temporal information with limited number of stored features in the memory, but also diversify states of a particular object to enhance the association performance. Our approach significantly improves over MOTRv2 in the DanceTrack test set, demonstrating a gain of 2.0% AssA score and 2.1% in IDF1 score.
@@ I Introduction
Historically, research efforts in Multiple Object Tracking (MOT) have primarily pursued two distinct directions: tracking-by-detection [sort, deepsort] methodologies and end-to-end models [centertrack, tracktor, memot, memotr]. Tracking-by-detection methods typically rely on object detectors to identify individual objects within each frame, followed by the application of filtering techniques like the Kalman filter to establish correspondence between bounding boxes across the video sequence [sort, ocsort, bytetrack]. Some innovations in this realm have focused on enhancing reidentification by leveraging deep features [deepsort, deepocsort, qdtrack]. In contrast, end-to-end models represent a more recent paradigm shift, consolidating both detection and association tasks within a unified deep-learning framework. Initially, earlier work often used CNN-based architecture, followed by the propagation of detected objects into subsequent frames [tracktor, centertrack]. However, with the introduction of the DETR (DEtection TRansformer) model [detr], which leverages the Transformer architecture [transformer] and stands as a state-of-the-art detector at that time, attention-based tracking approaches [trackformer, transtrack, motr], a subtype of end-to-end models, have emerged, demonstrating promising results in various tracking scenarios.
@@
Our proposed method offers a streamlined solution: a sparse memory that selectively stores critical features based on object motion and overlapping awareness. This approach leverages the intuition that an object’s degree of deformation correlates directly with its movement, thereby minimizing redundancy and enhancing long-temporal information. To validate our hypothesis, we incorporate our memory mechanism into the MOTRv2 [motrv2] model due to its inherent flexibility. MOTRv2 represents a novel hybrid approach, combining tracking-by-attention and tracking-by-detection methodologies, thus harnessing the strengths of robust detectors and a transformer-like architecture for tracking tasks. This architecture is adaptable across diverse scenarios due to its separation of detector and tracker components. By extending this flexibility, we introduce a novel training-free memory mechanism for object tracking, thereby enhancing the association performance of the model.
@@ II-A Tracking-by-detection
Previous methods in multi-object tracking (MOT) typically involved training a detector to locate objects in each frame individually, followed by linking these detections across the video sequence. One effective tracking method is shown in the SORT [sort] algorithm. SORT predicts where an object will go next using a Kalman filter and then uses the Hungarian algorithm to link these predictions with new detections. DeepSORT [deepsort] builds on SORT by adding more detailed object feature matching, which helps in identifying objects again after they disappear. Another notable approach, BYTETrack [bytetrack], handles situations where detections might have low confidence due to obstacles or blurriness. BYTETrack deals with these uncertain detections by associating them in a second pass, which reduces cases of wrongly identifying objects as absent. OC-SORT [ocsort] introduces an innovative update mechanism for the Kalman filter: when an object reappears after being lost, the filter’s parameters are adjusted based on the new observation. Deep OC-SORT [deepocsort] extends this idea by also updating deep features adaptively for object reidentification, guided by the confidence scores of detections.
@@ II-B End-to-end Approaches
End-to-end object tracking models utilize two main approaches: tracking-by-regression and tracking-by-attention. Tracking-by-regression propagates object detections from one frame to the next, bypassing traditional matching methods. For example, Tracktor [tracktor] uses a regression head to determine object existence in subsequent frames, while CenterTrack [centertrack] predicts detections based on object centers from previous frames. With the introduction of DETR [detr], a new approach emerged, integrating detection and tracking using Transformer architecture. TrackFormer [trackformer] builds on this, using an attention mechanism in the Transformer to treat objects from previous frames as queries for subsequent frames, enhancing tracking effectiveness. Similarly, MOTR [motr] extends this concept by aggregating track queries across multiple frames. TransTrack [transtrack] employs an IoU matching mechanism instead of a regression head for determining new object appearances or tracking previous ones.
II-C Hybrid Approaches and Memory-based Approaches
Although tracking-by-attention methods are effective, they struggle with adapting to diverse object classes, requiring simultaneous training for detection and tracking. MOTRv2 [motrv2] addresses this by combining tracking-by-detection and tracking-by-attention, using a pretrained detector and an attention mechanism to propagate objects across frames. Studies have shown that memory mechanisms enhance reidentification [memotr, memot] in MOT. Unlike SOT, which deals with fewer objects, MOT involves more objects, making memory management challenging. Therefore, instead of storing object features at each frame, we propose a novel approach for memory, which stores object sparsely.
@@ III Method
Figure 1: Our proposed method seamlessly integrates into the MOTRv2 [motrv2] model as a training-free module, preserving MOTRv2’s flexibility for various applications. As depicted in the figure, the pipeline takes proposals generated by YOLOX [yolox2021] and previously tracked objects (detailed in Section III-A) as input for MOTRv2. The Spatial-aware Sparse Memory (SASM), further explained in Figure 2, then processes MOTRv2’s outputs.
In this section, we present our method, which builds on the MOTRv2 [motrv2] model known for its flexibility and superior tracking capabilities. We enhance MOTRv2 by introducing a memory module to improve reidentification performance. As shown in Figure 1, our memory module filters the object embeddings produced by MOTRv2 in each frame to retain the most important features. These selected features are then propagated to the next frame to aid in object localization. The filtering process includes two modules: Spatial-Aware Sparse Memory (SASM) and Overlapping-Aware Feature Selector (OFS), detailed in Sections III-B and III-C respectively.
III-A Revisiting MOTRv2
Y Zhang et. al. propose integrating the YOLOX [yolox2021] object detector to provide detection priors for MOTR [motr]. By providing detection priors, YOLOX reduces the burden on the tracker, which would otherwise be responsible for handling both detection and tracking tasks.
III-A1 Proposal query generation
In this approach, for each frame $t$, YOLOX generates $N_{t}$ proposals $P$, each characterized by center coordinates $(x_{t},y_{t})$, height $h_{t}$, width $w_{t}$, and confidence score $s_{t}$. Initially, a query mechanism similar to the DAB-DETR [liu2022dabdetr] architecture is employed for the first frame The confidence scores are also integrated into the proposal queries $\mathbf{q}_{t}$ using sine-cosine positional encoding.
III-A2 Proposal propagation
@@
In subsequent frames, the object queries $\mathbf{q}_{t-1}^{tr}$ and predictions from the previous frame are propagated and concatenated with the proposals from the new frame $P_{t}$.
III-B Spatial-aware Sparse Memory
In this study, our objective is to optimize memory utilization in video processing by minimizing redundancy in object information storage. We propose a sparse storage strategy tailored to capture significant object deformations between frames, thereby avoiding unnecessary data retention during periods of minimal change. Acknowledging the diverse deformation characteristics across objects, a uniform storage approach proves inefficient. Consequently, we hypothesize a direct relationship between an object’s movement dynamics and its deformation profile in video sequences, motivating an adaptive memory management scheme.
Our approach leverages the displacement of each object’s centroid as a decisive criterion for selective memory storage. By dynamically adjusting storage intervals based on real-time deformation analysis, our method aims to enhance storage efficiency while minimizing redundancy information. Figure 2 delineates the architecture of our proposed memory management system, encompassing the following methodological steps:
Figure 2: The Spatial-aware Sparse Memory (SASM) takes the output of MOTRv2 as input, which includes both object queries and coordinates. From the object coordinates, the module filters out small-displacement objects to accumulate distance to the next frame, and stores only objects with large displacement, as described in Section III-B. The Overlapping-aware Feature Selector (OFS) plays a crucial role in selecting the best features of each instance since the last memory update, as described in Section III-C.
@@ III-B Spatial-aware Sparse Memory
    Accumulating Moving Distance: If the moving distance $d_{i,t}$ does not surpass $\epsilon$, it is accumulated with the moving distance of the object in the next frame $t+1$. This accumulation accounts for gradual movement across consecutive frames.
    $d_{i,t+1}=d_{i,t+1}+d_{i,t}$
    Otherwise, if the computed moving distance $d_{i,t}$ exceeds a predefined threshold $\epsilon$, indicating significant motion, the object’s features are stored in memory. This step ensures that only objects undergoing substantial movement are considered for memory storage. Additionally, with a higher $\epsilon$, longer temporal information of the object can be stored, but it may also lead to the omission of some important features.
@@
These steps collectively form a coherent framework for selective features stored in memory, leveraging object movement as a key criterion for memory allocation and ensuring the diversity of object state in object tracking across frames. As a result, this sparse memory strategy not only addresses the issue of bias towards the latest appearances of objects in feature representation but also facilitates the capture of longer temporal information by reducing redundancy when compared with dense memory.
III-C Overlapping-aware Feature Selector
Empirical observations reveal that object features are often incorrectly stored during occlusion or overlap, introducing noise. To address this, we propose selectively storing object features in memory.
A simplistic approach of waiting for unobstructed views leads to significant temporal gaps. Instead, we select the feature from the frame with the lowest Intersection over Union (IoU) with other objects since the last update. This minimizes erroneous features and ensures smoother continuity.
We define the stored feature $\mathbf{\hat{q}}_{i,t}$ at frame $t$ as:
@@ III-C Overlapping-aware Feature Selector
where $\mathbf{q}_{i,t}^{tr}$ is the feature of object $i$ at frame $t$, and $\mathbf{q}_{i,k}^{tr}$ is the feature at frame $k$ with the highest IoU. This selection process retains more reliable features, reducing noise and e
As a result, this selective feature storage strategy contributes to our tracking system’s overall efficiency and effectiveness, enabling more accurate and reliable object tracking over extended durations.
@@ IV-A Dataset and Metric
In this paper, we evaluate our methods on the DanceTrack [dancetrack] dataset. This dataset requires tracking multiple objects with uniform appearance and diverse motion.To ensure fair comparisons, we assess our methods using well-known metrics as employed in previous works [motr, motrv2]: HOTA [hota] to evaluate our method and analyze the contribution decomposed into detection accuracy (DetA) and association accuracy (AssA); MOTA [clearmot]; IDF1 [id-metrics] metrics.